A farmer has 2x chickens as ducks. After selling 166 chickens and 29 ducks he is left with 2x ducks as chickens. What is the original amount of the chickens
Answer by richwmiller(17219)

2d=c,
2*(c-166)=d-29
4d=d+303
3d=303
d=101
c=202
 202 chickens  101 ducks
-166           -29
  36            72